OpenStudy (anonymous):

4x - 2y = 6 2x + y = 5 Solve

OpenStudy (anonymous):

do you know how we would start

OpenStudy (anonymous):

yeah buy the 2x?

OpenStudy (anonymous):

i think elimination would be easiest to use

OpenStudy (anonymous):

do you know how to do elimination/substitution

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

\[4x-2y= 6 \]

OpenStudy (anonymous):

lets do this one and i try the other on my own?

OpenStudy (anonymous):

you have to have both in the first one lets do elimination 4x-2y=6 2x+y=5

OpenStudy (anonymous):

one of the y's are negative so lets look at the second equation to eliminate y we need them to be same number but opposite signs

OpenStudy (anonymous):

to get 2y we multiply the whole equation by 2 \[4x-2y=6\] \[4x+2y=10\]

OpenStudy (anonymous):

now we cancel out the y's\[4x=6\]\[4x=10\]

OpenStudy (anonymous):

combine both equations together \[8x=16\] solve for x

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2x?

OpenStudy (anonymous):

alright so x=2 now we substitute this in to one of the equations

OpenStudy (anonymous):

lets substitute it in for the second equation since y is by itself \[2(2)+y=5\]

OpenStudy (anonymous):

why do that?

OpenStudy (anonymous):

because we found x we substitute it in for 1 equation to solve for y

OpenStudy (anonymous):

ohhhh ok

OpenStudy (anonymous):

so can you solve for y?

OpenStudy (anonymous):

so what next

OpenStudy (anonymous):

multiply 2 and 2 together then subtract the product into 5

OpenStudy (anonymous):

y=?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

why would it be negative?

OpenStudy (anonymous):

you said to multiply 2 x 2 = 4 then subtract 5 right?

OpenStudy (anonymous):

no subtract 4 into 5

OpenStudy (anonymous):

5-4

OpenStudy (anonymous):

1 OMG

OpenStudy (anonymous):

yes lol x=2 y=1 (2,1)

OpenStudy (anonymous):

thanks :)